OpenStudy (anonymous):

Find two angles that satisfy the equation sin(3x + 9)° = cos(5x + 73)°. a.12;78 b.78;102 c.12;88

hartnn (hartnn):

do you know the relation, sin x = cos (90-x) ??

hartnn (hartnn):

you use the relation and you get 90-(3x+9) = 5x+73 can you find x from here ??

OpenStudy (anonymous):

The two angels are complimentery (sum is 90) so (3x+9)+(5x+73)=90 8x+82=90 8x=8 x=1 subsitut x answer is 12 &78